OpenStudy (anonymous):

Caitlin says the difference in their ages is 24 years. In two or more complete sentences, explain whether or not Caitlin is correct. In your explanation, include the equations and calculations necessary to support or contradict Caitlin's answer.

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

will medal

OpenStudy (anonymous):

@buckfever is dat all there must be more, a graph, perhaps a diagram, something.

OpenStudy (anonymous):

sorry here is the rest of the info

OpenStudy (anonymous):

The ratio of the ages of Mandy and Sandy is 2:5. After 8 years, their ages will be in the ratio 1:2. What is the difference between their present ages?

OpenStudy (anonymous):

24 is correct. The equations are shown below: {m/s = 2/5, (m + 8)/(s + 8) = 1/2}

OpenStudy (anonymous):

thx 4 da medal